OpenStudy (anonymous):

Which is larger 3/4 or 4/7 or are they equal?

OpenStudy (anonymous):

ur mom

OpenStudy (anonymous):

you have to find the least common denominator first

OpenStudy (zpupster):

find the common denominator: (3/4* 7/7) = ?? 4/7* 4/4) =??

OpenStudy (mokeira):

find the LCM and multiply both by the LCM...LCM=28 3/4(28)=21 4/7(28)=16 so 3/4 is larger